I need help with this
A Go-ped traveled 9 miles in 24 minutes.
How long would it take to go 27 miles?

: I need help with this
A Go-ped traveled 9 miles in 24 minutes.
How long would it take to go 27 miles?


Answer by stanbon(18792)

A Go-ped traveled 9 miles in 24 minutes.
How long would it take to go 27 miles?
--------------------------
Let time to go 27 miles be x-minutes.
Proportion:
x/27 = 24/9
9x = 27*24
x = 3*24
x = 72 minutes
=================
Cheers,
Stan H.